# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ 3.6 Example: retrieval-augmented generation
Note again that while we focus on this specific setting in our experiments, the algorithm and analysis in the following are actually applicable to generic settings of this RAG task.
3.6.1 Algorithm
@@
We note a few details about this algorithm. (1) For each chunk in Step 2, we prompt the LLM to retrieve relevant sentences, or return “None” if no relevant sentence is found in that chunk. Such “None” results will be excluded from Step 3 of the algorithm. (2) Unlike previous examples, the final aggregation step of this algorithm involves an LLM node, which adds to the cost metrics of the overall algorithm. (3) For simplicity, we assume that the number of needles (i.e. length of the passcode) and the length of each needle are both $O(1)$. For more general cases, the final aggregation step might benefit from further task decomposition. (4) We allow the algorithm to return a partial answer, by placing a special character in the digit(s) of the passcode that it is uncertain about.
3.6.2 Analysis
@@
Error metrics.
Our analysis of error metrics for this task is similar to that for the previous retrieval example. In particular, there are two possible failure modes in the retrieval step, and conclusions for the errors of the final solution returned by the overall algorithm are dependent on whether a Type-1 or Type-2 LLM is being used. For example, if a Type-1 LLM, which will not mistakenly retrieve irrelevant sentences from the input text, is used within the algorithm, then a smaller value of $m$ implies higher accuracy of retrieval in Step 2 of the algorithm, which further leads to lower error metrics for the solution returned by the final aggregation step.
Cost metrics.
For simplicity, let us assume that a Type-1 LLM is used within the overall algorithm. Consequently, in Step 2 of the algorithm, each LLM call has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(m)$ and $\mathcal{L}_{\textsf{dec}}=O(1)$, since only the relevant text within the chunk is retrieved. Moreover, among the $k$ LLM calls, only $O(1)$ of them return answers that are not “None”. By excluding the “None” results from the final aggregation step, the last LLM call has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(1)$ and $\mathcal{L}_{\textsf{dec}}=O(1)$. Putting things together, with a degree of parallelism $p$ and the number of chunks $k=\lceil 2n/m-1\rceil$, the cost metric $\mathcal{C}$ of the overall algorithm is bounded by
@@
3.6.3 Experiments
We empirically validate the performance of the proposed LLM-based algorithm with a Llama-3-70B model. Two error metrics are considered: the exact-match error taking value in $\{0,1\}$, and the fraction of incorrect digits, which takes value in $[0,1]$ and is always no larger than the exact-match error.
@@ 3.6.3 Experiments
In Figure 10(b), we fix $n=20000$ and vary the chunk size $m$. As was predicted by our analysis for a Type-1 LLM, a smaller value of $m$ implies lower error metrics, indicating the efficacy of task decomposition.
@@
Figure 10: Empirical results for RAG with Llama-3-70B.
3.7 Example: long-text summarization with chunking
For the final example of this section, we apply our analytical framework to the task of long-text summarization [KRA+22, CLGI24], where generating a summary for the long input text with a single LLM call is infeasible. Algorithms that process the long text to be summarized by chunks have been proposed, such as hierarchical merging [WOZ+21] and incremental updating [Ope24b]. The former first generates one summary for each chunk independently and then merge these intermediate summaries into the final one, while the latter maintains and updates a global summary as the chunks get processed one by one in order. Since objective and quantitative evaluation for the summarization task is known to be challenging [CLGI24] and remains an active research area, we focus our study on the cost metrics of both algorithms. Our analysis leads to interesting observations that are different from those in previous examples, primarily due to the various options of setting the number of generated tokens $\mathcal{L}_{\textsf{dec}}$ for each LLM call within the algorithms.
3.7.1 Processing chunks in parallel
@@
Let us consider the following algorithm that processes the chunks in parallel, which is a simplification of hierarchical merging [CLGI24] and visualized in Figure 2(a):
@@
Let us specify the hyperparameters $s_{1}$ and more generally $\{s_{i}\}$, assuming that the targeted length $s$ of the final summary has been determined. If the targeted information that we wish to be included in the final summary is distributed evenly across the input text, e.g. in a story with a linear narrative, then it is reasonable to set $s_{i}=s_{1}=s/k$ for all $i\in[k]$. In other scenarios where the targeted information is distributed unevenly (e.g. in query-focused summarization [LFXW24, VFK+22]), it is generally safer to set $s_{i}=s_{1}=s$ for all $i\in[k]$. The latency with parallelism degree $p$ in both cases is summarized as follows:
@@
3.7.2 Processing chunks sequentially
@@
Let us consider the following algorithm that processes the chunks sequentially, which is a simplification of incremental updating [CLGI24] and visualized in Figure 2(b):
@@
4 Case study: hierarchical decomposition
This section studies the pattern of hierarchical decomposition for LLM-based algorithms, where the original task is decomposed into multiple sub-tasks, and each of them can be further decomposed into more lower-level sub-tasks. This pattern is strictly more expressive than parallel decomposition studied in the previous section, and hence capable of solving more problems.
@@ 4 Case study: hierarchical decomposition
For concreteness, we investigate a challenging version of the RAG task studied in Section 3.6 that requires multi-hop reasoning [YQZ+18, LZLC24]. Recall that in the previous RAG example, given a question and multiple text chunks, the algorithm following the pattern of parallel decomposition will first retrieve useful sentences from each chunk, and then invoke one LLM call to answer the question based on the retrieved information. Such a method might not be feasible in more challenging scenarios, where the needles embedded in the haystack are logically related, and some of the needles are related to the targeted question only indirectly via connection to other needles. For example, suppose that the question is “what is the numeric value of A”, while the needles are “A = B”, “B = C”, and “C = 100”, located separately in different chunks. Retrieving needles from their corresponding chunks solely based on the targeted question will certainly fail in this case.
@@
To tackle this challenge, a natural idea is to extend the RAG algorithm considered in Section 3.6, allowing it to perform multiple rounds of iterative retrieval and reasoning. For each round, the algorithm performs retrieval based on the targeted question as well as additional information from previous rounds, followed by reasoning about the retrieved sentences and deciding whether the targeted question can be answered, or further retrieval and reasoning is needed. Similar approaches have been widely adopted in prior work, e.g. in the Selection-Inference framework [CSH23], in a RAG system with iterative follow-up questions for applications in medicine [XJW+24], or as a technique of extending the effective context window of a LLM-powered agent system [QT24]. The resulting algorithm for reasoning with iterative retrieval exhibits a hierarchical structure: the original task is decomposed into multiple sequential rounds, and each round is further decomposed into multiple steps of retrieval and reasoning.
@@
In the rest of this section, we first elaborate the concrete setup and algorithms for this case study, and then demonstrate our analysis of accuracy and efficiency based on the proposed analytical framework, as well as insights derived from it. Finally, numerical experiments are conducted to validate the efficacy and scalability of the considered algorithm, as well as our analysis and insights.
4.1 Concrete setup
@@
For concreteness, let us consider the task of finding the numeric value of a targeted variable embedded within a grid, similar to the setting considered in [YXLAZ24]. A visualization can be found in Figure 11. Configurations of the grid include its depth $d$, width $w$, and degree $g$, which control the difficulty of this task. More specifically, the grid consists of $d$ levels, each containing $w$ variables; each variable at a certain level is a function (e.g. addition, substraction, maximum or minimum) of $g$ variables at the next level, except for the variables at the final level, whose numeric values are given. Such information is provided in clues of the form “A2 = B1 + B3” or “C1 = 10”, each of which corresponds to one variable. As a result, the total number of clues is $d\times w$, and each clue has length $O(g)$. We refer to the clues that are necessary and sufficient for calculating the targeted variable as the needles, which constitute a directed acyclic graph (DAG) embedded within the grid. It is assumed that the level of reasoning required in this case study is non-trivial yet also simple enough, in the sense that one single LLM call with step-by-step reasoning is sufficient for answering the targeted question correctly if all needles are given a priori.
@@
(a) The “cyclic” version.
(b) The “parallel” version.
Figure 12: Algorithms for reasoning with iterative retrieval, which exhibit the pattern of hierarchical decomposition. While the number of rounds is assumed to be 2 in this visualization, it is actually determined adaptively by the algorithm itself at runtime. For clarity of visualization, some LLM or non-LLM nodes (cf. Figure 1) are merged into one, and an arrow from the “chunking” node to a dashed block means that each “retrieval” node within the dashed block takes the corresponding chunk as input.
4.2 Algorithm
@@
    Divide the input clues into $k$ chunks. In addition, initialize an empty list of references, i.e. clues retrieved by LLM calls, which will be maintained and updated throughout the algorithm.
@@
    • If the LLM cannot answer, and no additional clue has been retrieved in this round, then the algorithm terminates and fails [^6] This is due to the assumption that greedy decoding, which is deterministic, is adopted. In this case, running one more round of retrieval and reasoning will give the same result as that of the current round, which will be useless and wasteful. A potential improvement is to adaptively decrease (e.g. halve) the chunk size and continue the algorithm with more rounds; such improvements are not the focus of this work, and hence omitted here..
@@
We consider two options of prompting the LLM to reason about the references and answer the targeted question: answering directly, or thinking step by step before answering [KGR+22]. It is reasonable to expect that the latter will boost the accuracy while incurring higher costs, which will soon be investigated analytically and quantitatively in Section 4.3.
@@
With the “cyclic” option, the overall algorithm typically need fewer rounds to answer the targeted question correctly, since each LLM call for retrieval always leverages the most updated references. On the other hand, the “parallel” option can leverage parallelism of LLM calls for reducing the end-to-end latency of the overall algorithm. The comparison between these two options will soon be elaborated in Section 4.3.
Remark 3.
The algorithm under consideration exhibits the pattern of hierarchical decomposition, in that the overall task is decomposed into a sequence of rounds, and each round is further decomposed into multiple steps of retrieval and reasoning. Another feature of this algorithm is that its computational graph is constructed dynamically at runtime, since the number of rounds is determined adaptively by the algorithm itself.
4.3 Analysis
4.3.1 Error metrics
There are two major failure modes for this algorithm of reasoning with iterative retrieval:
  • Mistakes in reasoning and answering. For example, the LLM call responsible for reasoning and answering might mistakenly decide that it is not yet ready to answer the targeted question, or make mistakes when doing the arithmetic and thus return a wrong numeric value, even when all needles have been successfully retrieved. It is also possible that the LLM call mistakenly returns a numeric value (e.g. the value of a variable whose name is similar to that of the targeted variable), while the needles have not been fully retrieved yet. The probability of this failure mode is particularly related to whether the LLM call responsible for reasoning and answering is prompted to answer directly or think step by step before answering.
  • Failure to retrieve all needles. As long as the ground-truth DAG of needles has a limited size, the probability of this failure mode is largely determined by the choice of the chunk size. Since this has been thoroughly investigated in our previous retrieval (Section 3.5) and RAG (Section 3.6) examples, it will not be our focus in this section.
4.3.2 Cost metrics
@@
Recall from Eq. (2) that $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})$ stands for the cost of one LLM call with a prompt of length $\mathcal{L}_{\textsf{pre}}$ and generated text of length $\mathcal{L}_{\textsf{dec}}$. We also adopt some notations from Section 3.1 in our previous study of parallel decomposition: $n$ represents the total length (in tokens or characters) of the input clues, $m$ represents an upper bound for the chunk size, $k\lesssim n/m$ denotes the number of chunks, and $\mathcal{L}_{\textsf{sys}}$ denotes an upper bound for the length of the system prompt for each LLM call.
In addition, let $r$ denote the number of rounds, which is determined adaptively by the algorithm itself at runtime. Finally, let $\ell$ be the total length of the needles, i.e. clues that are necessary and sufficient for answering the targeted question correctly. The value of $\ell$ can depend on the depth $d$, width $w$ and degree $g$ of the grid in various ways:
@@
  • An LLM call responsible for reasoning and answering takes the references of length $O(\ell)$ as input, namely $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(\ell)$. As for the length of generated text $\mathcal{L}_{\textsf{dec}}$, it is reasonable to assume that $\mathcal{L}_{\textsf{dec}}\lesssim 1$ if the LLM is prompted to answer directly, and $\mathcal{L}_{\textsf{dec}}\lesssim\ell$ if it is prompted to do the calculation step by step before returning its final answer777 There might exist more precise characterizations of $\mathcal{L}_{\textsf{dec}}$ in this case, e.g. $\mathcal{L}_{\textsf{dec}}\lesssim\ell\times g$ if the LLM tends to take $g$ steps for calculating the sum of $g$ numbers. For simplicity, we mostly assume $g=O(1)$ in this case study, and thus stick to the assumption that $\mathcal{L}_{\textsf{dec}}\lesssim\ell$.. Consequently, one has
@@
Now we are ready to find out the cost of the overall algorithm with $r$ rounds of iterative retrieval (for $k$ chunks) and reasoning:
$\mathcal{C}$ $\leq r\times\mathcal{C}(\text{one round})$
$=r\times\Big{(}k\times\mathcal{C}(\text{retrieval})+\mathcal{C}(\text{reasoning})\Big{)}$
$\leq r\times\Big{(}k\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m+\ell,\ell)+\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)\Big{)}.$
In particular, this bound quantifies how the cost of LLM calls responsible for reasoning and answering, namely $r\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)$, only occupies a small fraction of the total cost of the overall algorithm.
@@
  • For the same task instance, the “parallel” version requires the same or a larger number of rounds, since in the “cyclic” version, each retrieval step always leverages the most updated references. For example, consider the case with degree $g=1$ and a chain of $d$ needles located separately in different chunks. Then, the number of rounds will be $r=d$ for the “parallel” version, but can be as small as $r=1$ for the “cyclic” version, if the needles happen to appear in the right order in the original clues.
@@
    $\mathcal{C}$ $\leq r\times\bigg{(}\lceil\frac{k}{p}\rceil\times\mathcal{C}(\text{retrieval})+\mathcal{C}(\text{reasoning})\bigg{)}$
    $\leq r\times\bigg{(}\lceil\frac{k}{p}\rceil\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m+\ell,\ell)+\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)\bigg{)}.$ (6)
@@
4.3.3 Insights
In sum, we derive two major insights from the above analysis of error and cost metrics:

### v3

@@ A.5 Example: retrieval-augmented generation
Note again that while we focus on this specific setting in our experiments, the algorithm and analysis in the following can be extended to generic settings of this RAG task.
A.5.1 Algorithm
@@
We note a few details about this algorithm. (1) For each chunk in Step 2, we prompt the LLM to retrieve relevant sentences, or return “None” if no relevant sentence is found in that chunk. Such “None” results will be excluded from Step 3 of the algorithm. (2) Unlike previous examples, the final aggregation step of this algorithm involves an LLM node, which adds to the cost metrics of the overall algorithm. (3) For simplicity, we assume that the number of needles (i.e., length of the passcode) and the length of each needle are both $O(1)$. For more general cases, the final aggregation step might benefit from further task decomposition. (4) We allow the algorithm to return a partial answer, by placing a special character in the digit(s) of the passcode that it is uncertain about.
A.5.2 Analysis
@@
Error analysis.
Our analysis of error metrics for this task is similar to that for the previous retrieval example. In particular, there are two possible failure modes in the retrieval step, and conclusions for the errors of the final solution returned by the overall algorithm are dependent on whether a Type-A or Type-B LLM is being used. For example, if a Type-A LLM, which will not mistakenly retrieve irrelevant sentences from the input text, is used within the algorithm, then a smaller value of $m$ implies higher accuracy of retrieval in Step 2 of the algorithm, which further leads to lower error metrics for the solution returned by the final aggregation step.
Cost analysis.
For simplicity, let us assume that a Type-A LLM is used within the overall algorithm. Consequently, in Step 2 of the algorithm, each LLM call has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(m)$ and $\mathcal{L}_{\textsf{dec}}=O(1)$, since only the relevant text within the chunk is retrieved. Moreover, among the $k$ LLM calls, only $O(1)$ of them return answers that are not “None”. By excluding the “None” results from the final aggregation step, the last LLM call has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(1)$ and $\mathcal{L}_{\textsf{dec}}=O(1)$. Putting things together, with a degree of parallelism $p$ and the number of chunks $k=\lceil 2n/m-1\rceil$, the end-to-end latency $\mathcal{C}$ of the overall algorithm is bounded by
@@
A.5.3 Experiments
We empirically validate the performance of the proposed LLM-based algorithm with a Llama-3-70B-Instruct model. Two error metrics are considered: the exact-match error taking value in $\{0,1\}$, and the fraction of incorrect digits, which takes value in $[0,1]$ and is always no larger than the exact-match error.
@@ A.5.3 Experiments
In Figure 15(b), we fix $n=20000$ and vary the chunk size $m$. As was predicted by our analysis for a Type-A LLM, a smaller value of $m$ implies lower error metrics, indicating the efficacy of task decomposition.
@@
Figure 15: Empirical results for RAG with Llama-3-70B-Instruct.
A.6 Example: long-text summarization with chunking
For the final example of parallel decomposition, we apply our analytical framework to the task of long-text summarization (Kryscinski et al., 2022; Chang et al., 2024), where generating a summary for the long input text with a single LLM call is infeasible. Algorithms that process the long text to be summarized by chunks have been proposed, such as hierarchical merging (Wu et al., 2021) and incremental updating (OpenAI, 2024b). The former first generates one summary for each chunk independently and then merge these intermediate summaries into the final one, while the latter maintains and updates a global summary as the chunks get processed one by one in order. Since objective and quantitative evaluation for the summarization task is known to be challenging (Chang et al., 2024) and remains an active research area, we focus our study on the cost metrics of both algorithms. Our analysis leads to interesting observations that are different from those in previous examples, primarily due to the various options of setting the number of generated tokens $\mathcal{L}_{\textsf{dec}}$ for each LLM call within the algorithms.
A.6.1 Processing chunks in parallel
@@
Let us consider the following algorithm that processes chunks in parallel, which is a simplification of hierarchical merging (Chang et al., 2024) and visualized in Figure 3(a):
@@
Let us specify the hyperparameters $s_{1}$ and more generally $\{s_{i}\}$, assuming that the targeted length $s$ of the final summary has been determined. If the targeted information that we wish to be included in the final summary is distributed evenly across the input text, e.g., in a story with a linear narrative, then it is reasonable to set $s_{i}=s_{1}=s/k$ for all $i\in[k]$. In other scenarios where the targeted information is distributed unevenly (e.g., in query-focused summarization (Laban et al., 2024; Vig et al., 2022)), it is generally safer to set $s_{i}=s_{1}=s$ for all $i\in[k]$. The latency with parallelism degree $p$ in both cases is summarized as follows:
@@
A.6.2 Processing chunks sequentially
@@
Let us consider the following algorithm that processes chunks sequentially, which is a simplification of incremental updating (Chang et al., 2024) and visualized in Figure 3(b):
@@
Appendix B Proof of Proposition 2
$\mathcal{E}(y(v))\leq\mathcal{E}_{v}=\sum_{w\in\mathcal{V}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$
$\mathcal{E}(y(u_{i}))\leq\sum_{w\in\mathcal{V}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w},\quad i\in[k].$
$\leq\mathcal{E}_{v}+S\times\sum_{i\in[k]}\sum_{w\in\mathcal{V}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+S\times\sum_{i\in[k]}\sum_{w\in\mathcal{V}\setminus\{v\}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+S\times\sum_{w\in\mathcal{V}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+\sum_{w\in\mathcal{V}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|+1}\times\mathcal{E}_{w}.$
Here in the third line, we replace the summation over $w\in\mathcal{V}$ with $w\in\mathcal{V}\setminus\{v\}$, since there is certainly no path from node $v$ to its predecessor node $u_{i}$. To move forward, notice that $\mathcal{P}(w\rightarrow v)$ is a union of $k$ disjoint sets, where the $i$-th set contains paths ending with the directed edge from $u_{i}$ to $v$:
With this, we can further simplify the previous upper bound:
$\mathcal{E}(y(v))$ $\leq\mathcal{E}_{v}+\sum_{w\in\mathcal{V}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|+1}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+\sum_{w\in\mathcal{V}\setminus\{v\}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\sum_{w\in\mathcal{V}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$
This concludes our proof of the proposition.
Appendix C Supplementary materials for Section 4.2
@@ Appendix B Proof of Proposition 2
This appendix includes full details of our case study, presented in Section 4.2, on iterative retrieval and reasoning. For concreteness, we investigate a challenging version of the RAG task studied in Section A.5 that requires multi-hop reasoning (Yang et al., 2018; Li et al., 2024a). Recall that in the previous RAG example, given a question and multiple text chunks, the algorithm (with the pattern of parallel decomposition) will first retrieve useful sentences from each chunk, and then invoke one LLM call to answer the question based on the retrieved information. Such a method might not be feasible in challenging scenarios where the needles embedded in the haystack are logically related, and some of the needles are related to the targeted question only indirectly via connection to other needles. For example, suppose that the question is “what is the numeric value of A”, while the needles are “A = B”, “B = C”, and “C = 100”, located separately in different chunks. Retrieving needles from their corresponding chunks solely based on the targeted question will certainly fail in this case.
@@
To tackle this challenge, a natural idea is to extend the RAG algorithm considered in Section A.5, allowing it to perform multiple rounds of iterative retrieval and reasoning. For each round, the algorithm performs retrieval based on the targeted question as well as additional information collected from previous rounds, followed by reasoning about the retrieved sentences and deciding whether the targeted question can be answered, or further retrieval and reasoning is needed. Similar approaches have been widely adopted in prior work, e.g., in the Selection-Inference framework (Creswell et al., 2023), in a RAG system with iterative follow-up questions for applications in medicine (Xiong et al., 2024), as a technique of extending the effective context window of a LLM-powered agent system (Qwen-Team, 2024), among others (Yue et al., 2024). The resulting algorithm for reasoning with iterative retrieval exhibits a hierarchical structure: the original task is decomposed into multiple sequential rounds, and each round is further decomposed into multiple steps of retrieval and reasoning.
@@
In the rest of this section, we first elaborate the concrete setup and algorithms for this case study, and then demonstrate our analysis of accuracy and efficiency based on the proposed analytical framework, as well as insights derived from it. Finally, numerical experiments are conducted to validate the efficacy of the considered algorithm, as well as the derived analysis and insights.
C.1 Concrete setup
@@
For concreteness, let us consider the task of finding the numeric value of a targeted variable embedded within a grid, similar to the setting considered in (Ye et al., 2024). A visualization can be found in Figure 16. Configurations of the grid include its depth $d$, width $w$, and degree $g$, which control the difficulty of this task. More specifically, the grid consists of $d$ levels, each containing $w$ variables; each variable at a certain level is a function (e.g., addition, substraction, maximum or minimum) of $g$ variables at the next level, except for the variables at the final level, whose numeric values are given. Such information is provided in clues of the form “A2 = B1 + B3” or “C1 = 10”, each of which corresponds to one variable. As a result, the total number of clues is $d\times w$, and each clue has length $O(g)$. We refer to the clues that are necessary and sufficient for calculating the targeted variable as the needles, which constitute a directed acyclic graph (DAG) embedded within the grid. It is assumed that the level of reasoning required in this case study is non-trivial yet also simple enough, in the sense that one single LLM call with step-by-step reasoning is sufficient for answering the targeted question correctly if all needles are given a priori.
@@
C.2 Algorithm
@@
    Divide the input clues into $k$ chunks. In addition, initialize an empty list of references, i.e., clues retrieved by LLM calls, which will be maintained and updated throughout the algorithm.
@@
    • If the LLM cannot answer, and no additional clue has been retrieved in this round, then the algorithm terminates and fails [^5] This is due to the assumption that greedy decoding, which is deterministic, is adopted. In this case, running one more round of retrieval and reasoning will give the same result as that of the current round, which will be useless and wasteful. A potential improvement is to adaptively decrease (e.g., halve) the chunk size and continue the algorithm with more rounds; such improvements are not the focus of this work, and hence omitted here..
@@
We consider two options of prompting the LLM to reason about the references and answer the targeted question: answering directly, or thinking step by step before answering (Kojima et al., 2022). It is reasonable to expect that the latter will boost the accuracy while incurring higher costs, which will soon be investigated analytically and quantitatively in Section C.3.
@@
With the “cyclic” option, the overall algorithm typically need fewer rounds to answer the targeted question correctly, since each LLM call for retrieval always leverages the most updated references. On the other hand, the “parallel” option can leverage parallelism of LLM calls for reducing the end-to-end latency of the overall algorithm. The comparison between these two options will soon be elaborated in Section C.3.
C.3 Analysis
C.3.1 Error analysis
There exist two major failure modes for this algorithm:
  1. 1.
    Mistakes in reasoning and answering. For example, the LLM call responsible for reasoning and answering might mistakenly decide that it is not yet ready to answer the targeted question, or make mistakes when doing the arithmetic and thus return a wrong numeric value, even when all needles have been successfully retrieved. It is also possible that the LLM call mistakenly returns a numeric value (e.g., the value of a variable whose name is similar to that of the targeted variable), while the needles have not been fully retrieved yet. The probability of this failure mode is particularly related to whether the LLM call responsible for reasoning and answering is prompted to answer directly or think step by step before answering.
  2. 2.
    Failure to retrieve all needles. As long as the ground-truth DAG of needles has a limited size, the probability of this failure mode is largely determined by the choice of the chunk size. Since this has been thoroughly investigated in our previous retrieval (Section A.4) and RAG (Section A.5) examples, it will not be our focus in this section.
C.3.2 Cost analysis
@@
Recall from Eq. (2) that $\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})=\mathcal{C}_{\textsf{pre}}(\mathcal{L}_{\textsf{pre}})+\mathcal{C}_{\textsf{dec}}(\mathcal{L}_{\textsf{pre}},\mathcal{L}_{\textsf{dec}})$ stands for the cost of one LLM call with a prompt of length $\mathcal{L}_{\textsf{pre}}$ and generated text of length $\mathcal{L}_{\textsf{dec}}$. We also adopt some notations from Section 3 in our previous study of parallel decomposition: $n$ represents the total length (in tokens or characters) of the input clues, $m$ represents an upper bound for the chunk size, $k\lesssim n/m$ denotes the number of chunks, and $\mathcal{L}_{\textsf{sys}}$ denotes an upper bound for the length of the system prompt for each LLM call.
In addition, let $r_{\textsf{cyc}}$ (resp. $r_{\textsf{par}}$) denote the number of rounds determined adaptively by the “cyclic” (resp. “parallel”) algorithm itself at runtime. Finally, let $\ell$ be the total length of the needles, i.e., clues that are necessary and sufficient for answering the targeted question correctly. The value of $\ell$ can depend on the depth $d$, width $w$ and degree $g$ of the grid in various ways:
@@
  • An LLM call responsible for reasoning and answering takes the references of length $O(\ell)$ as input, namely $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(\ell)$. As for the length of generated text $\mathcal{L}_{\textsf{dec}}$, it is reasonable to assume that $\mathcal{L}_{\textsf{dec}}\lesssim 1$ if the LLM is prompted to answer directly, and $\mathcal{L}_{\textsf{dec}}\lesssim\ell$ if it is prompted to do the calculation step by step before returning its final answer [^6] There might exist more precise characterizations of $\mathcal{L}_{\textsf{dec}}$ in this case, e.g., $\mathcal{L}_{\textsf{dec}}\lesssim\ell\times g$ if the LLM tends to take $g$ steps for calculating the sum of $g$ numbers. For simplicity, we mostly assume $g=O(1)$ in this case study, and thus stick to the assumption that $\mathcal{L}_{\textsf{dec}}\lesssim\ell$.. Consequently, one has
@@
Now we are ready to find out the cost of the overall algorithm with $r_{\textsf{cyc}}$ rounds of iterative retrieval (for $k$ chunks) and reasoning:
$\mathcal{C}$ $\leq r_{\textsf{cyc}}\times\mathcal{C}(\text{one round})$ (7)
$=r_{\textsf{cyc}}\times\Big(k\times\mathcal{C}(\text{retrieval})+\mathcal{C}(\text{reasoning})\Big)$ (8)
$\leq r_{\textsf{cyc}}\times\Big(k\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m+\ell,\ell)+\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)\Big).$ (9)
In particular, this bound quantifies how the cost of LLM calls responsible for reasoning and answering, namely $r_{\textsf{cyc}}\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)$, only occupies a small fraction of the total cost of the overall algorithm.
@@
  • For the same task instance, the “parallel” version requires the same or a larger number of rounds ($r_{\textsf{par}}\geq r_{\textsf{cyc}}$), since in the “cyclic” version, each retrieval step always leverages the most updated references. For example, consider the case with degree $g=1$ and a chain of $d$ needles located separately in different chunks. Then, the number of rounds will be $r_{\textsf{par}}=d$ for the “parallel” version, but can be as small as $r_{\textsf{cyc}}=1$ for the “cyclic” version, if the needles happen to appear in the right order in the original clues.
@@
    $\mathcal{C}$ $\leq r_{\textsf{par}}\times\bigg(\lceil\frac{k}{p}\rceil\times\mathcal{C}(\text{retrieval})+\mathcal{C}(\text{reasoning})\bigg)$
    $\leq r_{\textsf{par}}\times\bigg(\lceil\frac{k}{p}\rceil\times\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+m+\ell,\ell)+\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+\ell,1\text{ or }\ell)\bigg).$ (10)
@@
C.3.3 Insights
In sum, we derive two major insights from the above analysis: